### v1

# How to Bridge the Sim-to-Real Gap in Digital Twin-Aided Telecommunication Networks

Abstract
Training effective artificial intelligence models for telecommunications is challenging due to the scarcity of deployment-specific data. Real data collection is expensive, and available datasets often fail to capture the unique operational conditions and contextual variability of the network environment. Digital twinning provides a potential solution to this problem, as simulators tailored to the current network deployment can generate site-specific data to augment the available training datasets. However, there is a need to develop solutions to bridge the inherent simulation-to-reality (sim-to-real) gap between synthetic and real-world data. This paper reviews recent advances on two complementary strategies: 1) the calibration of digital twins (DTs) through real-world measurements, and 2) the use of sim-to-real gap-aware training strategies to robustly handle residual discrepancies between digital twin-generated and real data. For the latter, we evaluate two conceptually distinct methods that model the sim-to-real gap either at the level of the environment via Bayesian learning or at the level of the training loss via prediction-powered inference.
I Introduction
Driven by the continued growth of computing resources and training datasets, artificial intelligence (AI) research is widely considered to be in the scaling era, which is focused on the development of general-purpose models that exhibit emergent capabilities. While this trend has yielded impressive results for many tasks, particularly in the domain of language modeling, it poses unique challenges when applied to engineering domains such as telecommunication networks.
One major concern is the energy cost: training and deploying large-scale AI models in data centers consumes significant amounts of power, raising sustainability and operational efficiency issues. Another critical challenge is the highly contextual nature of telecommunication data. The performance of wireless networks is highly sensitive to specific deployment factors, including the propagation environment, network topology, and traffic patterns. These factors vary widely across different settings, making it impractical to collect sufficient real-world data to train robust AI models for every possible context.
A promising approach to address data scarcity in telecommunications is the use of digital twins (DTs)—high-fidelity simulators or emulators of network environments that can efficiently generate synthetic data tailored to specific deployments. By training AI models on a combination of real and synthetic data, it becomes possible to improve generalization performance compared to training solely on the limited real-world data typically available for a given deployment. Examples of DTs include NVIDIA’s Sionna library, which offers an efficient simulator of the propagation environment via ray tracing [1], with recent extensions also incorporating networking aspects [2]. DTs may also encompass real-world elements such as software-defined radios [3].
Figure 1: (left) Illustration of the sim-to-real gap between (top) a real-world urban wireless deployment and (bottom) a ray tracing-based digital twin (DT). The DT model may yield incorrect channel impulse responses (CIRs) due to a mismatch between the real geometry and material parameters and the ones assumed by the simulator. (right) When using the DT-predicted CIRs to train a beamforming model, the sim-to-real gap may cause the selection of a sub-optimal beam.
However, a key challenge when training on a mix of real and synthetic data is that the DT may not perfectly reflect the real data distribution. There is typically a non-negligible simulation-to-reality (sim-to-real) gap due to modeling simplifications or uncertain environment parameters. The left part of Fig. 1 illustrates this gap for a ray tracing DT engine (bottom) that fails to correctly simulate the true wireless propagation conditions of the real urban environment (top) due to its simplified virtual geometry and its lack of knowledge of the real electromagnetic parameters of the objects in the scene. If an AI model is trained naively on synthetic data produced by such a DT, these discrepancies can cause degraded performance when the model is deployed into the real-world system. In the right part of the example in Fig. 1, a beamforming model trained on synthetic data (bottom) may fail to select the best beam (top).
How can we bridge the sim-to-real gap to ensure that AI models trained with DTs remain reliable in real deployments? As illustrated in Fig. 2(a), this article explores two complementary strategies:
$\bullet$ Calibration of DTs (Sec. II): We first discuss techniques for calibrating DTs using real-world measurements, with the goal of directly improving simulator fidelity and reducing bias in synthetic data generation. For example, in the setting illustrated in Fig. 1, calibration methods may improve the description of the scene geometry and/or the estimate of the material parameters in the scene by matching real-world observations with synthetic data (see Fig. 2(a)).
$\bullet$ Robust AI training via DTs (Sec. III and IV): Calibration methods may reduce the sim-to-real gap, but they cannot fully remove it owing to the inherent bias entailed by the use of limited-complexity models and simulators. Therefore, even when using a calibrated DT, it is important to design training methods that explicitly account for the residual sim-to-real gap in the synthetic data generated by the DT. These techniques fall into two distinct, and complementary, categories:
  • $\circ$
    Modeling the sim-to-real gap on the environment (Sec. III): A first approach is to explicitly model the sim-to-real gap directly at the level of the DT model. Accordingly, the underlying DT model is treated as a random quantity, on which the DT refines its prior knowledge by using the available data. In the ray tracing example, the geometry and/or the electromagnetic properties of the objects in the scene may be modeled as random variables, whose uncertainty describes the residual ambiguity about the mismatch between real and simulated environments (see Fig. 2(b)). This probabilistic modeling, formalized by Bayesian inference, makes it possible to generate synthetic data from multiple plausible DT models, thus accounting for the sim-to-real gap resulting from prior knowledge and from limited data.
  • $\circ$
    Modeling the sim-to-real gap on the training objective (Sec. IV): Synthetic data produced by a DT are typically used to train AI models, e.g., the beamforming mapping in Fig. 1. In these settings, training is often done by minimizing a training loss evaluated using both real and synthetic data. In light of this, a complementary approach to devising reliable AI training strategies in the presence of DTs is to mitigate the sim-to-real gap directly at the level of the training loss. These techniques seek to reduce the DT bias under sim‑to‑real gaps by incorporating available real‑world observations. A prominent approach to achieve this is prediction-powered inference (PPI) [4] (see Fig. 2(c)).
(a) Digital twin calibration
(c) Modeling the sim-to-real gap on the training objective
Figure 2: As detailed in this paper, the sim-to-real gap can be partially bridged via three main complementary strategies: (a) Digital twin calibration (Sec. II), which uses real-world measurements to directly improve the adherence of the DT to the real world; (b)-(c) Robust AI training, which accounts for the residual sim-to-real gap by modeling the uncertainty on the true environment via Bayesian inference (panel (b), Sec. III), or by correcting the bias caused by the sim-to-real gap on the training objective (panel (c), Sec. IV). (Dashed lines indicate DT-generated synthetic data, while solid lines represent real-world measurements.)
(b) Modeling the sim-to-real gap on the environment
II Calibrating Digital Twins
Emerging DTs of telecommunication networks often include models of the wireless propagation environment using ray tracing [1]. In a ray tracing simulator, the environment is defined by a three-dimensional geometry along with a vector of material parameters describing the reflection, refraction, and diffraction characteristics of surfaces in the scene. The simulator models the propagation of radio waves by tracing the paths connecting transmitters to receivers, keeping track of the interactions of each path with the objects in the environment. The multi-path parameters returned by the simulator can then be leveraged to produce a simulated channel response or to estimate the received power at given locations (see Fig. 1).
In practice, however, a sim-to-real gap is inevitable: the specified geometry of the scene might miss objects or report inaccurate placements or dimensions, while material parameters may be unknown or wrongly specified. These mismatches lead to biases in the synthetic data generated by the DT. Therefore, it is important to reduce the sim-to-real gap caused by model misspecification by leveraging available real-world data to calibrate the DT.
II-A Material and Parameter Calibration via Differentiable Simulation
One way to calibrate a DT is to learn the unknown material parameters by comparing simulation outputs to real measurements. To elaborate, we denote as $\rho$ the material parameters for the objects present in the scene. These include permittivities, conductivities, and scattering patterns. In practice, it is also possible to define an implicit neural representation (INR) that associates each position in the scene with a vector $\rho$ [1].
To support calibration, the network operator deploys a few test transmitters and receivers in the real network, and records the channel measurements $H$. Given a vector of model parameters $\rho$, one can then run the ray tracing simulator to yield a prediction $\hat{H}(\rho)$ of the channel measurements. Using this information, calibration can be formulated as an inverse problem: find the parameters $\hat{\rho}$ that cause the simulated measurement $\hat{H}(\hat{\rho})$ to best match the real observations $H$. This criterion is averaged over the available data set in order to obtain statistically sound estimates of the parameters $\rho$.
Recent research has shown that this inverse problem can be efficiently addressed by leveraging differentiable ray tracing software, and applying gradient-based optimization techniques to adjust the material parameters [1]. This approach treats the DT as a trainable model mapping material parameters to multi-path parameters, making it possible to differentiate through the ray tracing engine. This, in turn, supports the use of gradient descent to infer material properties $\rho$ that minimize the error between simulated and measured channel responses.
II-B Phase Error-Aware Calibration
Calibrating material parameters is useful, but it cannot fully resolve biases caused by a misspecified geometry of the objects in the scene. For instance, as seen in Fig. 2(a), minor errors in building shapes or additional objects not present in the DT can cause significant phase errors in the simulated propagation paths. In fact, even a small displacement of an object—of the order of a fraction of the wavelength—can cause the phase of a reflected path to shift, leading to constructive or destructive interference patterns that do not align with reality [5]. Owing to these unaccounted phase errors, traditional calibration methods that only tweak material values $\rho$ may fail to produce an accurate simulator.
To tackle this issue, a promising approach is to devise calibration schemes that are aware of the presence of phase errors caused by a sim-to-real gap in the geometry of the environment. The core idea behind phase error-aware calibration, as proposed in [5], is to iteratively estimate and correct phase errors in the simulated channel. In this method, the DT’s prediction is augmented with per-path phase error terms that account for unmodeled geometry nuances. An expectation-maximization (EM) algorithm is then used: in the E-step, the algorithm infers the likely phase errors for each propagation path by comparing the simulation with real measurements; while the M-step updates the simulator parameters $\rho$ to better fit the observed data, effectively absorbing the phase errors into the model.
Fig. 3 shows the performance of different calibration techniques for a canyon scenario with line-of-sight (LoS) blockage under geometry mismatch at the DT [5]. The calibration procedures are conducted under multiple levels of available bandwidth for channel measurements, and the results quantify the relative error between the measured power and the power predicted by the calibrated DT. We compare the described phase error-aware approach with the following alternative methods: (i) phase error-oblivious calibration, which disregards phase errors due to geometry mismatch when matching simulated and measured channel responses; and (ii) uniform phase error calibration, which assumes the phases of the predicted paths to be uniformly distributed when matching delay-angle projections of the predicted and measured channel powers during calibration. All methods tune the material parameters $\rho$ via gradient-based optimization [1] (see Sec. II-A).
By disregarding phase errors, phase-oblivious calibration is seen to be unable to provide satisfactory results across all bandwidth levels. Uniform phase error calibration yields improved results as the bandwidth levels approach the delay-resolution needed to separate the contributions of the interfering paths. Finally, phase error-aware calibration can dramatically reduce the power prediction error as compared to the two baseline methods, even at lower bandwidth levels. This demonstrates that explicitly accounting for local phase error during calibration can significantly improve the reliability of the ray tracer’s estimates. This advantage is especially pronounced in scenarios with a rich multipath structure, in which small geometry differences would otherwise lead to large interference errors.
Figure 3: Relative power prediction errors of ray tracing-based DTs calibrated using phase error-oblivious (orange), uniform phase error (green), and phase error-aware (blue) calibration methods, where calibration is conducted using channel measurements at different bandwidths (figure adapted from [5]).
III Modeling the Sim-to-Real Gap on the Environment
Through calibration, the synthetic data produced by the DT becomes a much closer proxy for real data. Calibration thus attacks the sim-to-real gap at its source, reducing the bias in any model trained on the DT. Nonetheless, in practice, there will always remain a residual uncertainty or bias, contributing to a non-negligible sim-to-real gap. For example, a ray tracing simulator will inevitably leverage simplified geometric and electromagnetic configurations. For this reason, it is also important to design techniques that make the AI training process itself robust to any residual sim-to-real gap.
Rather than further refining the simulator, which may not be feasible, the goal of the next two sections is to robustly train AI models using synthetic data generated by a DT. A principled way to account for the sim-to-real gap during training is to explicitly model uncertainty about the true environment at the DT itself before generating synthetic data. This way, rather than assuming the digital replica of the environment provided by the DT to be ideal, we treat the parameters that define the DT —such as the material or geometric parameters discussed in Sec. II—as random variables within a Bayesian perspective.
Figure 4: Training a multiple access protocol using synthetic data from a digital twin (DT) modeling the access channel: (left) a Bayesian formulation of DT learning assigns different configurations (interference levels) of the DT to a probability distribution that depends on prior knowledge and available data. In contrast, a frequentist approach would only select the highest-probability model. (center) While the frequentist DT trains the policy inside the single selected model, the Bayesian DT generates data from multiple models, weighting the importance of multiple plausible DT models during policy optimization. (right) When calibration data is in limited supply, explicitly modeling the sim-to-real gap via uncertainty-aware Bayesian methods can significantly improve the performance of models trained using synthetic data from the DT (adapted from [6]).
In this Bayesian framework, as illustrated in Fig. 2(b), one maintains a probability distribution over the possible configurations of the DT, reflecting the epistemic uncertainty on the environment under the limited available data. Initially, this uncertainty is encoded in a prior distribution, which expresses our knowledge—or lack thereof—before observing any real-world data. As we collect real measurements, the distribution is updated to a posterior distribution, thereby refining the DT’s understanding of the environment and modeling any residual sim-to-real gap. Instead of relying on a single “best-guess” DT model for synthetic data generation, with Bayesian learning one can generate synthetic data from an ensemble of DT models sampled from the posterior distribution, enabling a more robust learning of AI policies. By sampling models from the posterior distribution, during data generation, the weight given to each DT model depends on the belief assigned by the system to the given model based on the available information.
As an exemplifying use case, following [6], we now focus on a system in which multiple devices coordinate to access a shared wireless medium using a protocol learned via multi-agent reinforcement learning (MARL). The MARL protocol is trained using synthetic data from a DT. If we were to use a single DT model to simulate the environment, as done in a conventional frequentist implementation, the resulting policy might perform well under that specific model, but it could fail to generalize to the real-world system due to the real-to-sim gap. In contrast, the Bayesian approach samples multiple plausible environments from the posterior distribution, training the MARL policy across a number of different DT models. This leads to a policy that is robust under the sim-to-real gap modeled at the DT, achieving better generalization to the true environment.
Fig. 4 elaborates on the operation and on the advantages of the Bayesian methodology for the given multiple access setting. The left panel illustrates a posterior distribution over different DT models that assume varying levels of interference, which result in distinct likelihoods to the number of received packets given the number of concurrently transmitted packets. The central panel contrasts the training processes of different methodologies. The conventional frequentist learning approach relies on a single maximum a posteriori model, while the Bayesian learning framework integrates across a range of likely models, weighting each model according to its posterior probability. As shown in the right panel, this yields superior performance for the policy trained using synthetic data, especially in data-scarce regimes where the model uncertainty about the environment dynamics is high.
Bayesian DT training can be implemented through various practical methods. Examples include Monte Carlo dropout, deep ensembles, and variational inference techniques that approximate the posterior distribution [7, Ch. 12]. It should be emphasized that, while these approaches improve robustness and reliability, they do come at the cost of increased computational complexity: sampling multiple environments, training across a population of models, and maintaining distributions over parameters may introduce significant overhead. Nonetheless, in many critical applications where reliability and safety are paramount, this additional cost often offers a worthwhile trade-off.
IV Modeling the Sim-to-Real Gap on the Training Objective
Another approach to address the residual sim-to-real gap is to leave the DT as is, but modify the training algorithm to compensate for the DT’s bias. A recently developed framework called PPI [4, 8] offers a powerful tool in this regard. PPI originates from the idea of doubly robust estimation in statistics, and it supports the use of abundant biased data from a simulator, together with a small amount of unbiased data from the real world, to produce unbiased learning outcomes.
IV-A Semi-Supervised Learning via Prediction-Powered Inference
PPI supports the definition of a semi-supervised training methodology that leverages both real and synthetic data. To elaborate, suppose we want to train an AI model for a certain task using mainly synthetic data generated by the DT. Assume also that we also have a limited set of real data for the same task, but not enough to train a high-capacity model from scratch. In classical transfer learning, one might pre-train on synthetic data and then fine-tune on real data [7, 9].
In contrast, as illustrated in Fig. 2(c), PPI uses the real data not to train the model directly, but to estimate the bias of a training loss estimate based on synthetic data. It then adjusts the model training objective to remove this bias. Essentially, the real data provides a yardstick for how far off the synthetic training loss is from the real training loss. PPI is appealing for telecommunication AI problems because obtaining even a small validation set of real data is often feasible through field tests or pilot deployments, whereas gathering a comprehensive training set is not [10].
The corrected training objective produced by PPI is guaranteed to be unbiased with respect to the real data distribution. In this regard, a key advantage is that the real data are used very efficiently, as they are not wasted on brute-force training, but rather on pinpointing the bias. This approach reduces reliance on real data by leveraging less expensive synthetic samples, enabling models to perform well in practice even when only a small fraction, e.g. $1$-$10\%$, of the training data is real [10].
(a) Cross-PPI
$2\cdot 10^{-2}$$4\cdot 10^{-2}$$6\cdot 10^{-2}$$8\cdot 10^{-2}$0.10.10.10.13.53.53.53.544444.54.54.54.555555.55.55.55.5Proportion of real dataChannel capacity (bps/Hz)ERMP-ERMPPICross-PPI
(b) Cross-PPI on beam alignment via CKM
Figure 5: (a) Illustration of Cross-PPI [8, 10]: The real data is divided into $K$ folds, and $K$ calibration procedures are conducted, each using all real data except one fold, held out for bias estimation. The calibration bias is estimated and removed from the synthetic data loss (as shown in Fig. 2(c)), obtaining the Cross-PPI loss. (b) Channel capacity as a function of the proportion of real data for Cross-PPI, PPI [4], empirical risk minimization (ERM), and pseudo-empirical risk minimization (P-ERM). As detailed in [10], given real data combined with synthetic samples, a classification model is trained to map UE-location to the optimal downlink beam index. The performance is measured in terms of achievable downlink channel capacity.
That said, PPI in its basic form has two important limitations when it comes to applications to telecommunication systems:
$\bullet$ Data efficiency: The requirement to hold out a portion of real data exclusively for bias correction means that those data cannot be used to calibrate the DT. If real data are extremely scarce, this may be a significant drawback.
$\bullet$ Context independence: The bias correction in vanilla PPI is typically a single global adjustment. In other words, it assumes the simulator’s bias is the same in all situations. In telecommunication contexts, the accuracy of a DT may vary with the scenario. For example, the DT might be very accurate for urban microcell cases, but less so for rural macrocells, or accurate for certain ranges of signal-to-noise ratio but not others [11]. A single bias term would struggle to correct all contexts equally.
IV-B Cross-Validation-Based Prediction-Powered Inference
Cross-PPI improves data efficiency by effectively recycling real data during calibration while still using it for bias estimation. As depicted in Fig. 5(a), this is achieved through a cross-validation-style procedure in which the real data is partitioned into $K$ folds, and synthetic data is calibrated using $K$ calibration procedures, where each procedure uses all real data except one fold, reserved for bias correction [8, 10]. In such a way, all real data is used for both calibration and bias correction.
To illustrate the advantages of Cross-PPI, we consider the problem of beam alignment via channel knowledge maps (CKMs) in mmWave massive MIMO systems [12]. A CKM is a site‑specific database of channel characteristics associated with the locations of both the transmitter and the receiver [12]. Recently, this concept has been leveraged to mitigate training overhead, offering a promising new approach for channel estimation and beam alignment [12].
The objective is to train a classification model that maps the UE location to the optimal downlink beam selected from a predefined finite codebook. Fig. 5(b) shows the performance of the trained model in terms of achievable downlink channel capacity as a function of the fraction of real samples. Three baselines are considered: empirical risk minimization (ERM), which ignores the synthetic data; pseudo-empirical risk minimization (P-ERM), which treats the synthetic data as ground‑truth samples; and PPI [4]. The results indicate that Cross‑PPI achieves substantial improvements compared to all other methods. Notably, P-ERM does not yield any clear benefits over ERM, which relies exclusively on the real data. Although PPI shows some improvements over the baselines, its performance remains lower than that of Cross‑PPI.
(a) Real data
(b) Synthetic data
(c) ERM
(d) PPI
(e) CPPI
Figure 6: (a) Heatmap for the ground-truth optimal azimuth angles required to serve a user in the given position from the BS placed at the black cross mark. The layouts of the buildings are depicted as empty white regions. (b) Optimal azimuth angles as prediced by a DT. (c)-(e) Heatmap for the pointing error obtained by a model trained using real and synthetic data via (c) empirical risk minimization (ERM), (d) PPI [13], and (e) context-aware PPI (CPPI) [11].
IV-C Context-Aware Prediction-Powered Inference
Context-aware PPI tackles the lack of contextuality of conventional PPI by allowing the bias correction applied to the training loss to vary with context. The core idea is to partition the synthetic data, and corresponding real validation data, as a function of relevant context features, and then perform bias estimation separately for each context. For instance, one can assign different correction factors for different ranges of a radio channel condition metric or for different cell sectors in a network. By doing so, the model treats synthetic data from contexts where the twin is reliable with more trust, and discounts or heavily corrects synthetic data in contexts where the twin is less accurate [11].
In Fig. 6, following [11], we consider an outdoor urban environment with a single base station (BS), where the propagation paths between the BS and the devices are generated using ray‑tracing as in [12]. Each device position is associated with a ground‑truth azimuth and elevation angle of departure (AoD), which is determined by selecting the propagation path with the highest received power. The objective is to train a mapping that outputs an estimate of the optimal AoD at each receive location, given a real dataset and a much larger synthetic dataset. Each data point is augmented with a binary context variable indicating if the device is within the LoS of the BS.
Panels (a) and (b) present geographic heatmaps of beamforming directions: panel (a) shows the ground truth obtained from real data, while panel (b) displays the prediction made by a DT. Notably, substantial discrepancies emerge in the non-line-of-sight (NLoS) region, where the DT fails to capture the true beamforming pattern, yielding a large sim-to-real gap. Panels (c)–(e) quantify the beam-pointing error for three training schemes: ERM, which relies solely on real data, PPI [13, 4], and context-aware PPI (CPPI) [11], respectively. As seen, PPI performs poorly in the NLoS locations, exhibiting even worse performance than ERM. In contrast, CPPI leverages context-dependent weightings of synthetic samples, thus achieving uniformly lower errors across the entire region.
V Conclusion
As AI continues to scale, its promise for revolutionizing network automation and optimization is tempered by the realities of deployment in complex, data-scarce telecommunication environments. DTs have emerged as a vital tool to generate the data needed to train and test AI solutions virtually. By calibrating DTs with real-world data, and by explicitly modeling residual sim-to-real gaps through Bayesian and PPI techniques, engineers can enhance the performance and reliability of DT-aided telecommunication networks.
The techniques discussed in this paper are key to designing AI models that harness the full power of modern machine learning-based simulation tools, while maintaining reliability in the face of real-world network complexities. However, several challenges remain. Network non-stationarity demands adaptive calibration, where approaches like meta-learning frameworks [14] may be highly valuable. Furthermore, distributional shifts can erode uncertainty estimates, making the integration of techniques such as adaptive sampling [15] crucial to maintaining the robustness of PPI. Finally, generating synthetic samples within the DT, whether via ray tracing or ML‑based emulation, incurs tangible costs. Carefully balancing the volume of synthetic data against the budgetary constraints and the targeted model performance is essential for successful deployment.